# Charges, complex structures, and perturbations of instantons

### v3

@@ 2.3. Conformal transformations
where $\Omega$ is a positive scalar field. The metric on self-dual 2-forms (2.2) transforms as $\hat{I}_{ABCD}=\Omega^{2}I_{ABCD}$. Since the Weyl spinor is conformally invariant, $\hat{\Psi}_{ABCD}=\Psi_{ABCD}$, then $\hat{\Psi}_{AB}{}^{CD}=\Omega^{-2}\Psi_{AB}{}^{CD}$. From (2.9), the eigenvalues of $\Psi_{AB}{}^{CD}$ transform as $\hat{\lambda}_{i}=\Omega^{-2}\lambda_{i}$. If $\phi^{i}_{AB}$ is an orthonormal basis of eigenspinors of $\Psi_{AB}{}^{CD}$, then a basis of eigenspinors of $\hat{\Psi}_{AB}{}^{CD}$ which is orthonormal w.r.t. $\hat{I}_{ABCD}$ is $\hat{\phi}^{i}_{AB}=\Omega\phi^{i}_{AB}$. Formulas (2.9), (2.10), (2.11) remain valid for the conformally transformed variables, after putting hats over all quantities. Consider one of the real self-dual 2-forms in (2.1), say $Z_{ab}\equiv Z^{3}_{ab}$. Then $J^{a}{}_{b}=\sqrt{2}Z_{bc}g^{ca}$ is an almost-complex structure compatible with $g_{ab}$. Under a conformal change (2.12), $\hat{Z}_{ab}=\Omega^{2}Z_{ab}$, so $J^{a}{}_{b}$ is conformally invariant. The Lee form $f=f_{a}{\rm d}{x}^{a}$ (see e.g. [36]) is defined by
@@
(In indices, $\partial_{[a}Z_{bc]}=-2f_{[a}Z_{bc]}$.) If $(M,g_{ab})$ is Einstein [^3] All of our results for Einstein manifolds apply in particular in the Ricci-flat case. We shall not mention this explicitly in the following. and Hermitian w.r.t. $J^{a}{}_{b}$, with $W^{+}_{abcd}\neq 0$, then the Goldberg-Sachs theorem implies that $Z_{ab}$ is an eigenform of $W^{+}{}_{ab}{}^{cd}$:
@@ 3.1. Charges for Hermitian instantons
and $\sigma^{0}_{A^{\prime}},\sigma^{1}_{A^{\prime}}$ are defined in (A.7).
@@
The proof of (3.3) is straightforward if one uses a `conformal-GHP connection' [7]: this is an operator $\mathcal{C}_{AA^{\prime}}$ which generalizes the ordinary GHP connection [50, §4.12] to be invariant under combined conformal transformations and rescalings of the spinors $o_{A},\iota_{A}$, by taking into account both the conformal and the GHP weights of the objects it acts on. The necessary background is given in appendix A. Since $\zeta_{AB}$ has conformal weight $w=1$, then $\zeta^{AB}$ has w=1w=-1, and since $\Psi_{ABCD}$ has $w=0$, then $\psi_{AB}$ has w=1w=-1. Thus:
@@
where in the first line we used (A.5), in the third we used (A.6), and in the fourth we used (A.7) and the definition (3.5). ∎
@@
so the term with $S$ in (3.6) vanishes. Suppose that, in addition, $(M,g_{ab})$ is Hermitian, with complex structure $J^{a}{}_{b}=\sqrt{2}Z_{bc}g^{ca}$ (so that the principal spinors are $o_{A},\iota_{A}$, recall (3.1) and (2.1c)). Then from appendix A, we have $\sigma^{0}_{A^{\prime}}=0$ and the term with K*K in (3.6) vanishes. Recalling that for an Einstein-Hermitian geometry with $\Psi_{2}\neq 0$, the Lee form is (2.15), the result (3.6) reduces to
@@
One way of interpreting this identity is in terms of spin-lowering via Killing spinors [51, §6.4]. We have $\Psi_{2}^{-1/3}\Sigma_{ab}\propto\Psi_{ABCD}K^{CD}\epsilon_{A^{\prime}B^{\prime}}$, where $K_{AB}=\Psi_{2}^{-1/3}o_{(A}\iota_{B)}$ is a valence-2 Killing spinor, $\nabla_{A^{\prime}(A}K_{BC)}=0$ (cf. [51, Eq. (6.7.15)]). Then (3.9) is the statement that
@@
is an anti-self-dual Maxwell field. As explained by Penrose and Rindler in [51, §6.4], this construction can be applied to linearized gravity on Minkowski space-time by replacing $\Psi_{ABCD}$ with a spin 2 zero-rest-mass field. Then one can associate a charge to this field by integrating (3.10) on a closed 2-surface, where, since Minkowski space-time is topologically trivial, the 2-surface must surround a region of matter sources in order for the charge to be non-vanishing.
For our Einstein Hermitian geometry $(M,g_{ab})$, an analogous non-vanishing charge requires the existence of non-contractible 2-surfaces in $M$. Additionally, we can also interpret (3.9) in terms of complex structures: using (3.7) and (2.14), eq. (3.9) is a reflection of the fact that $(M,g_{ab})$ is conformally Kähler, with the Kähler form given by (2.16). Combining the two interpretations above, we can then define a charge associated to every 2-cycle $S$ in $M$, by formula (1.4). Note that this charge has a sort of combination of geometrical and topological nature, since it depends on both the curvature of $g_{ab}$ and the existence of non-trivial 2-cycles in $M$.
3.2. Perturbations and closed 2-forms
Consider a smooth one-parameter family of metrics $g_{ab}(s)$ in $M$. A tensor field $T(s)$ constructed from the metric has an expansion
$T(s)=T+(\delta{T})\,s+\frac{1}{2!}(\delta^{2}{T})\,s^{2}+O(s^{3}),$ (3.11)
where $T,\delta{T},\delta^{2}{T},$... are tensor fields in $M$,
@@ 3.2. Perturbations and closed 2-forms
We say that $T=T(s)\big|_{s=0}$ is the ``background'' value of $T(s)$, and $\delta{T}$ is the linearization of $T(s)$ at $T$. We recall that, under an infinitesimal diffeomorphism generated by an arbitrary vector field $\xi^{a}$, the linearization $\delta{T}$ transforms as
@@
which in this context is referred to as a gauge transformation. We say that the linearization $\delta{T}$ is gauge invariant if it is invariant under (3.13) for all $\xi^{a}$. Note that, in particular, this holds if the background value of $T(s)$ vanishes, i.e. $T=0$.
The Levi-Civita connection of $g_{ab}(s)$ is denoted $\nabla^{(s)}_{a}$, and its relation to the background Levi-Civita connection is (see [54, Eq. (7.5.7)])
$\nabla^{(s)}_{a}v_{b}=\nabla_{a}v_{b}-C_{ab}{}^{c}(s)\,v_{c}$ (3.14)
for all $v_{a}$, where
$C_{ab}{}^{c}(s)=\frac{1}{2}g^{cd}(s)\left[\nabla_{a}g_{bd}(s)+\nabla_{b}g_{ad}(s)-\nabla_{d}g_{ab}(s)\right]$ (3.15)
and $\nabla_{a}=\nabla_{a}^{(0)}$. Note that $C_{ab}{}^{c}(s)=s\,\delta{C}_{ab}{}^{c}+O(s^{2})$.
@@ 4. Examples
In this section we shall see several examples of the charges (1.4), by computing the conformal Kähler structures, the 2-cycles, and then the periods (1.4), for both compact and non-compact instantons.
A non-trivial issue is how to find the independent 2-cycles, for which we need the second homology group $H_{2}(M)$. To do this, we shall use the fact that all of the examples are toric. This implies that one can apply the rod structure formalism as in [19], which allows to determine the topological properties of the instanton. For example, for a rod structure with $n$ turning points, the Euler characteristic $\chi(M)$ is equal to $n$. In the non-compact case, we see from [47, Theorem 1] that $H_{2}(M)\cong\mathbb{Z}^{n-1}$. In the compact case, we can use the fact that $\chi(M)$ is also the alternating sum of the Betti numbers $b_{i}(M)$ to deduce the number of independent 2-cycles.
@@ 4.1. Ambi-Kähler examples
define integrable almost-complex structures with opposite orientation. The Kähler forms are then given by $\hat{\omega}^{\pm}=(\Psi_{2}^{\pm})^{2/3}\omega^{\pm}$. The topology is $\mathbb{R}^{2}\times S^{2}$, and the rod structure has two turning points, two semi-infinite rods and one finite rod [19, §4.4]. Since there are two turning points, we have $H_{2}(M)\cong\mathbb{Z}$ [47, Thm. 1], so there is only one independent 2-cycle $S$, corresponding to the bolt at $r=r_{+}$ of the Killing field $\partial_{\tau}+\Omega_{E}\partial_{\phi}$, where $r_{+}=m+\sqrt{m^{2}+a^{2}}$ and $\Omega_{E}=a/(2mr_{+})$. The associated charges (1.4) are
@@
The self-dual Taub-NUT instanton is obtained from (4.5) when $m=n$. It is ALF, and the topology is $S^{4}\setminus\{\rm pt.\}\cong\mathbb{R}^{4}$. From (4.6), we see that $\Psi^{-}_{2}=0$ and $\Psi^{+}_{2}=2n/(r+n)^{3}$, that is $W^{-}_{abcd}=0$ and $W^{+}_{abcd}\neq 0$, hence the space is hyper-Kähler with respect to one orientation, i.e. there is a 2-sphere of Kähler structures, and type D with respect to the other. The 2-forms (4.7) are however not Kähler; the geometry is not only hyper-Kähler but also strictly conformally Kähler w.r.t. both orientations [8]. The rod structure has one turning point and two semi-infinite rods [19, §4.2]. However, since the manifold is topologically trivial, there are no 2-cycles; thus the parameter $n$ does not correspond to a charge integral like (1.4). For a different construction involving 3-cycles, see [17].
@@
The Taub-bolt instanton is obtained from (4.5) when $m=\frac{5}{4}n$. It is ALF, and the topology is $\mathbb{CP}^{2}\setminus\{{\rm pt.}\}$. We have $\Psi_{2}^{-}=n/[4(r-n)^{3}]$, $\Psi_{2}^{+}=9n/[4(r+n)^{3}]$, so both $W^{\pm}_{abcd}$ remain type D. The rod structure has two turning points, two semi-infinite rods and one finite rod [19, §4.7]. As in Kerr we have two turning points, so $H_{2}(M)\cong\mathbb{Z}$ [47, Thm. 1] and there is one independent 2-cycle, which we denote by $S_{2n}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=2n$. We then find:
@@
where $f(r)=1-(a/r)^{4}$, and $a$ is a real constant. The topology is $T^{*}\mathbb{CP}^{1}$, and the asymptotics is ALE. The metric is Ricci-flat, the self-dual Weyl tensor is $W^{+}_{abcd}=0$, and the anti-self-dual Weyl tensor $W^{-}_{abcd}$ is type D  [^4] As noted by Gibbons and Pope [35], the Eguchi-Hanson metric (4.9) can be obtained from (4.5) by setting $m=n+a^{4}/(128n^{3})$, $r=m+\tilde{r}^{2}/(8n)$, and taking the limit $n\to\infty$ (and then redefining $\tilde{r}\to r$), see [35, Eqs. (4.15)-(4.16)]. This procedure leads to (4.10) (i.e., $W^{-}_{abcd}\neq 0$ and $W^{+}_{abcd}=0$, which is the opposite to self-dual Taub-NUT).:
@@
give integrable almost-complex structures $J^{\pm}$ with opposite orientation, and that ${\rm d}\omega^{+}=0$, so $J^{+}$ is Kähler. However ${\rm d}\omega^{-}\neq 0$, and this gives the conformally Kähler form $\hat{\omega}^{-}=(\Psi_{2}^{-})^{2/3}\omega^{-}$ needed to define the charge (1.4). The rod structure has two turning points, two semi-infinite rods and one finite rod [19, §4.5]. As in previous cases, two turning points give $H_{2}(M)\cong\mathbb{Z}$ [47, Thm. 1] so there is one independent 2-cycle, which we denote by $S_{a}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=a$. We have

### v4

@@ 3.1. Conformal geometry
where $\Omega$ is a positive scalar field. Consider one of the real self-dual 2-forms in (2.1), say $Z_{ab}\equiv Z^{3}_{ab}$. Then $J^{a}{}_{b}=\sqrt{2}Z_{bc}g^{ca}$ is an almost-complex structure compatible with $g_{ab}$. Under a conformal change (3.1), $\hat{Z}_{ab}=\Omega^{2}Z_{ab}$, so $J^{a}{}_{b}$ is conformally invariant. The Lee form $f=f_{a}{\rm d}{x}^{a}$ (see e.g. [28]) is defined by
@@
(In indices, $\partial_{[a}Z_{bc]}=-2f_{[a}Z_{bc]}$.) The principal spinors of $Z_{ab}$ are $o_{A},\iota_{A}$. Note that $J^{a}{}_{b}$ is invariant under the ``GHP'' transformation
$o_{A}\to zo_{A},\qquad\iota_{A}\to z^{-1}\iota_{A},$ (3.3)
$g_{ab}\to\Omega^{2}g_{ab},\qquad o_{A}\to\Omega^{1/2}o_{A},\qquad\iota_{A}\to\Omega^{1/2}\iota_{A}.$ (3.4)
A spinor/tensor field $\varphi^{A...A^{\prime}...}_{B...B^{\prime}...}$ which transforms under (3.3)-(3.4) as $\varphi^{A...A^{\prime}...}_{B...B^{\prime}...}\to\Omega^{w}z^{p}\varphi^{A...A^{\prime}...}_{B...B^{\prime}...}$ is said to have conformal weight $w$ and GHP weight $p$. These `weighted' fields $\varphi^{A...A^{\prime}...}_{B...B^{\prime}...}$ are to be thought of as sections of vector bundles over $M$. A linear connection $\mathcal{C}_{AA^{\prime}}$ on these bundles was constructed in [7]. If $\phi_{AB...K}$ is a totally symmetric spinor field with conformal weight $w$ and GHP weight $p$, we have:
$\mathcal{C}_{AA^{\prime}}\phi_{BC...L}={}$ $\nabla_{AA^{\prime}}\phi_{BC...L}+wf_{AA^{\prime}}\phi_{BC...L}+pP_{AA^{\prime}}\phi_{BC...L}$ (3.5)
where the 1-forms $f_{a}$ and $P_{a}$ are
$f_{a}=-\frac{1}{2}J^{c}{}_{b}\nabla_{c}J^{b}{}_{a},\qquad P_{a}=\frac{i}{2}(\Gamma_{a}{}^{2}{}_{1}-J^{b}{}_{a}f_{b}),$ (3.6)
and $\Gamma_{a}{}^{2}{}_{1}$ was defined in (2.5c). The 1-form $f_{a}$ is the Lee form, cf. (3.2).
Remark 3.1.
We shall later need the following facts:
  1. (1)
    If $\phi_{AB...K}$ has w=1w=-1, $p=0$, then
    $\mathcal{C}_{A^{\prime}}{}^{B}\phi_{AB...K}=\nabla_{A^{\prime}}{}^{B}\phi_{AB...K},$ (3.7)
    whereas for $w=0$, $p=0$:
    $\mathcal{C}_{A^{\prime}}{}^{B}\phi_{AB...K}=\nabla_{A^{\prime}}{}^{B}\phi_{AB...K}+f_{A^{\prime}}{}^{B}\phi_{AB...K}.$ (3.8)
  2. (2)
    $o_{A}$ and $\iota_{A}$ have weights $(w=\frac{1}{2},p=1)$ and $(w=\frac{1}{2},p=-1)$ respectively, so a calculation using (3.5) gives:
    $\mathcal{C}_{AA^{\prime}}o_{B}=\sigma^{0}_{A^{\prime}}\iota_{A}\iota_{B},\qquad\sigma^{0}_{A^{\prime}}=o^{A}o^{B}\nabla_{AA^{\prime}}o_{B},$ (3.9a)
    $\mathcal{C}_{AA^{\prime}}\iota_{B}=\sigma^{1}_{A^{\prime}}o_{A}o_{B},\qquad\sigma^{1}_{A^{\prime}}=\iota^{A}\iota^{B}\nabla_{AA^{\prime}}\iota_{B}.$ (3.9b)
    Note that $\sigma^{1}_{A^{\prime}}=(\sigma^{0}_{A^{\prime}})^{\dagger}$.
  3. (3)
    $(M,g_{ab})$ is Hermitian with respect to $J^{a}{}_{b}$ if and only if $\sigma^{0}_{A^{\prime}}=0$.
If $(M,g_{ab})$ is Einstein [^2] All of our results for Einstein manifolds apply in particular in the Ricci-flat case. We shall not mention this explicitly in the following. and Hermitian with respect to $J^{a}{}_{b}$, with $W^{+}_{abcd}\neq 0$, then the Goldberg-Sachs theorem implies that $Z_{ab}$ is an eigenform of $W^{+}{}_{ab}{}^{cd}$:
@@ 3.2. Charges for Hermitian instantons
and $\sigma^{0}_{A^{\prime}},\sigma^{1}_{A^{\prime}}$ are defined in (3.9).
@@
The proof of (3.15) is straightforward if one uses the conformal-GHP connection from section 3.1. Since $\zeta_{AB}$ has conformal weight $w=1$, then $\zeta^{AB}$ has w=1w=-1, and since $\Psi_{ABCD}$ has $w=0$, then $\psi_{AB}$ has w=1w=-1. Thus:
@@
where in the first line we used (3.7), in the third we used (3.8), and in the fourth we used (3.9) and the definition (3.17). ∎
@@
so the term with $S$ in (3.18) vanishes. Suppose that, in addition, $(M,g_{ab})$ is Hermitian, with complex structure $J^{a}{}_{b}=\sqrt{2}Z_{bc}g^{ca}$ (so that the principal spinors are $o_{A},\iota_{A}$, recall (3.13) and (2.1c)). Then from Remark 3.1, we have $\sigma^{0}_{A^{\prime}}=0$ and the term with K*K in (3.18) vanishes. Recalling that for an Einstein-Hermitian geometry with $\Psi_{2}\neq 0$, the Lee form is (3.11), the result (3.18) reduces to
@@
One way of interpreting this identity is in terms of spin-lowering via Killing spinors [41, §6.4]. We have $\Psi_{2}^{-1/3}\Sigma_{ab}\propto\Psi_{ABCD}K^{CD}\epsilon_{A^{\prime}B^{\prime}}$, where $K_{AB}=\Psi_{2}^{-1/3}o_{(A}\iota_{B)}$ is a valence-2 Killing spinor, $\nabla_{A^{\prime}(A}K_{BC)}=0$ (cf. [41, Eq. (6.7.15)]). Then (3.21) is the statement that
@@
is an anti-self-dual Maxwell field. As explained by Penrose and Rindler in [41, §6.4], this construction can be applied to linearized gravity on Minkowski space-time by replacing $\Psi_{ABCD}$ with a spin 2 zero-rest-mass field. Then one can associate a charge to this field by integrating (3.22) on a closed 2-surface, where, since Minkowski space-time is topologically trivial, the 2-surface must surround a region of matter sources in order for the charge to be non-vanishing.
For our Einstein Hermitian geometry $(M,g_{ab})$, an analogous non-vanishing charge requires the existence of non-contractible 2-surfaces in $M$. Additionally, we can also interpret (3.21) in terms of complex structures: using (3.19) and (3.10), eq. (3.21) is a reflection of the fact that $(M,g_{ab})$ is conformally Kähler, with the Kähler form given by (3.12). Combining the two interpretations above, we can then define a charge associated to every 2-cycle $S$ in $M$, by the formula (1.3).
3.3. Linearized charges
Given a smooth one-parameter family of tensor fields $T(s)$ in $M$, we define the $k$-jets
@@ 3.3. Linearized charges
The zeroth jet $T=T(0)$ is the background value of $T(s)$, and the first jet $\delta{T}$ is the linearization of $T(s)$ at $T$. We recall that, under an infinitesimal diffeomorphism generated by an arbitrary vector field $\xi^{a}$, the linearization $\delta{T}$ transforms as
@@
which in this context is referred to as a gauge transformation. We say that the linearization $\delta{T}$ is gauge invariant if it is invariant under (3.24) for all $\xi^{a}$. This holds, in particular, if $T=0$.
@@ 4. Examples
In this section we shall analyze several examples of the charges (1.3), by computing the conformal Kähler structures, the 2-cycles, and then the periods (1.3), for both compact and non-compact instantons. For each of these examples, Theorem 3.4 implies that generic gravitational perturbations also have a quasi-locally conserved charge, given by (1.4).
A non-trivial issue is how to find the independent 2-cycles, for which we need the second homology group $H_{2}(M)$. To do this, we shall use the fact that all of the examples are toric. This implies that one can apply the rod structure formalism as in [15], which allows to determine the topological properties of the instanton. For example, for a rod structure with $n$ turning points, the Euler characteristic $\chi(M)$ is equal to $n$. In the non-compact case, we see from [37, Theorem 1] that $H_{2}(M)\cong\mathbb{Z}^{n-1}$. In the compact case, we can use the fact that $\chi(M)$ is also the alternating sum of the Betti numbers $b_{i}(M)$ to deduce the number of independent 2-cycles.
@@ 4.1. Ambi-Kähler examples
define integrable almost-complex structures with opposite orientation. The Kähler forms are then given by $\hat{\omega}^{\pm}=(\Psi_{2}^{\pm})^{2/3}\omega^{\pm}$. The topology is $\mathbb{R}^{2}\times S^{2}$, and the rod structure has two turning points, two semi-infinite rods and one finite rod [15, §4.4]. Since there are two turning points, we have $H_{2}(M)\cong\mathbb{Z}$ [37, Thm. 1], so there is only one independent 2-cycle $S$, corresponding to the bolt at $r=r_{+}$ of the Killing field $\partial_{\tau}+\Omega_{E}\partial_{\phi}$, where $r_{+}=m+\sqrt{m^{2}+a^{2}}$ and $\Omega_{E}=a/(2mr_{+})$. The associated charges (1.3) are
@@
The self-dual Taub-NUT instanton is obtained from (4.5) when $m=n$. It is ALF, and the topology is $S^{4}\setminus\{\rm pt.\}\cong\mathbb{R}^{4}$. From (4.6), we see that $\Psi^{-}_{2}=0$ and $\Psi^{+}_{2}=2n/(r+n)^{3}$, that is $W^{-}_{abcd}=0$ and $W^{+}_{abcd}\neq 0$, hence the space is hyper-Kähler with respect to one orientation, i.e. there is a 2-sphere of Kähler structures, and type D with respect to the other. The 2-forms (4.7) are however not Kähler; the geometry is not only hyper-Kähler but also strictly conformally Kähler w.r.t. both orientations [8]. The rod structure has one turning point and two semi-infinite rods [15, §4.2]. However, since the manifold is topologically trivial, there are no 2-cycles; thus the parameter $n$ does not correspond to a charge integral like (1.3). For a different construction involving 3-cycles, see [13].
@@
The Taub-bolt instanton is obtained from (4.5) when $m=\frac{5}{4}n$. It is ALF, and the topology is $\mathbb{CP}^{2}\setminus\{{\rm pt.}\}$. We have $\Psi_{2}^{-}=n/[4(r-n)^{3}]$, $\Psi_{2}^{+}=9n/[4(r+n)^{3}]$, so both $W^{\pm}_{abcd}$ remain type D. The rod structure has two turning points, two semi-infinite rods and one finite rod [15, §4.7]. As in Kerr we have two turning points, so $H_{2}(M)\cong\mathbb{Z}$ [37, Thm. 1] and there is one independent 2-cycle, which we denote by $S_{2n}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=2n$. We then find:
@@
where $f(r)=1-(a/r)^{4}$, and $a$ is a real constant. The topology is $T^{*}\mathbb{CP}^{1}$, and the asymptotics is ALE. The metric is Ricci-flat, the self-dual Weyl tensor is $W^{+}_{abcd}=0$, and the anti-self-dual Weyl tensor $W^{-}_{abcd}$ is type D  [^3] As noted by Gibbons and Pope [27], the Eguchi-Hanson metric (4.9) can be obtained from (4.5) by setting $m=n+a^{4}/(128n^{3})$, $r=m+\tilde{r}^{2}/(8n)$, and taking the limit $n\to\infty$ (and then redefining $\tilde{r}\to r$), see [27, Eqs. (4.15)-(4.16)]. This procedure leads to (4.10) (i.e., $W^{-}_{abcd}\neq 0$ and $W^{+}_{abcd}=0$, which is the opposite to self-dual Taub-NUT).:
@@
give integrable almost-complex structures $J^{\pm}$ with opposite orientation, and that ${\rm d}\omega^{+}=0$, so $J^{+}$ is Kähler. However ${\rm d}\omega^{-}\neq 0$, and this gives the conformally Kähler form $\hat{\omega}^{-}=(\Psi_{2}^{-})^{2/3}\omega^{-}$ needed to define the charge (1.3). The rod structure has two turning points, two semi-infinite rods and one finite rod [15, §4.5]. As in previous cases, two turning points give $H_{2}(M)\cong\mathbb{Z}$ [37, Thm. 1] so there is one independent 2-cycle, which we denote by $S_{a}$ and corresponds to the bolt of $\partial_{\psi}$ at $r=a$. We have